OpenStudy (anonymous):

find the distance between points B(3,-2) and D(3,0) to the nearest tenth

OpenStudy (anonymous):

\[distance =\sqrt{(x-x1)^2+(y-y1)^2}...........d=\sqrt{(3-3)^2+(-2-0)^2}=2\]